OpenStudy (anonymous):

Four particles A, B, C and D of mass 2 kg, 5 kg, 6 kg and 3 kg respectively are rigidly joined together by light rods to form a rectangle ABCD where AB=2a and BC=4a. Find the moment of inertia of this system of particles about an axis along: a) AB b) BC c) AC d) through the midpoints of AB and CD

sam (.sam.):

\[I=mr ^{2}\]

OpenStudy (anonymous):

thank you..but, still not helping..i got that a lot in my note and book..

sam (.sam.):

AB, I=(6+3)(4a)^2 using formula mr^2

OpenStudy (anonymous):

there u are..does for D, I=(mL^2)/3..then (3+2)(2a)^2/3..=(20a^2)/3..does the question mean that the answer will be unknown?

OpenStudy (anonymous):

 as moment of inertia for a particle is I=mr 2 about ab means look at other masses other than aand b